OpenStudy (anonymous):

graph by plotting points. 2x=-y+3 (-1, ), (0, ) complete the ordered pairs

OpenStudy (anonymous):

2(-1) = y+3 -- solve for y 2(0) = y+3

OpenStudy (anonymous):

ok so how do i find the y

OpenStudy (anonymous):

2(-1) = y+3 -- simplify by multiplying 2 and -1 -2 = y+3 -- subtract 3 to both sides -3 -3 -------------------- -5 = y

OpenStudy (anonymous):

solve for the second part

OpenStudy (anonymous):

should have been 2(-1) = - y +3 then y = 5

OpenStudy (anonymous):

oops, didn't see the negative sign